# Day-Ahead Electricity Price Forecasting Using Merit-Order Curves Time Series

### v1

@@ Abstract
We introduce a general, simple, and computationally efficient framework for predicting day-ahead supply and demand merit-order curves, from which both point and probabilistic electricity price forecasts can be derived. Specifically, we leverage functional principal component analysis to efficiently represent a pair of supply and demand curves in a low-dimensional vector space and employ regularized vector autoregressive models for their prediction. We conduct a rigorous empirical comparison of price forecasting performance between the proposed curve-based model, i.e., derived from predicted merit-order curves, and state-of-the-art price-based models that directly forecast the clearing price, using data from the Italian day-ahead market over the 2023–2024 period. Our results show that the proposed curve-based approach significantly improves both point and probabilistic price forecasting accuracy relative to price-based approaches, with average gains of approximately 5%, and improvements of up to 10% during mid-day hours, when prices occasionally drop due to high renewable generation and low demand.
@@
electricity price forecasting , day-ahead market , merit-order curves , functional data analysis[mox]organization=MOX, Department of Mathematics, Politecnico di Milano,addressline=Piazza Leonardo da Vinci 32, city=Milano, postcode=20133, state=MI, country=Italy
@@ 1.1 Context
Because of this single-price property, a large body of research has focused on forecasting the day-ahead clearing price itself, a discipline known as electricity price forecasting (EPF). The consensus is that data-driven approaches based on historical prices tend to outperform fundamental (or structural) ones which try to model exact market mechanisms, especially for short-term hourly price forecasting (weron_electricity_2014). In particular, a highly competitive day-ahead EPF benchmark based on regularized autoregressive models and feedforward neural networks was consolidated by lago_forecasting_2021. This framework is still state-of-the-art and has been widely employed in recent studies (olivares_neural_2023; lipiecki_postprocessing_2024; oconnor_review_2025; cerasa_enhancing_2025).
@@
Nevertheless, the problem of forecasting the entire merit-order curves underlying price formation is gaining momentum in the academic community (petropoulos_forecasting_2022). The seminal work of ziel_electricity_2016 demonstrated the capacity of ”data-driven fundamental” approaches based on historical auction data and their potential for generating competitive clearing price forecasts. Forecasting entire curves offers several advantages. For participants with market power, it enables precise quantification of how marginal changes in offers may influence the clearing price and allows realistic market simulations from which optimal bidding strategies can be derived (grigoletto_supply_2013). For price-takers and other actors exposed to the day-ahead price, curve-based models offers a more comprehensive view of market results, enhances the interpretability of price predictions, and may improve the quantification of price predictions’ uncertainty222This latter aspect is investigated in this paper.. Additionally, complex nonlinear dynamics observed at the price level may become more tractable when modeled at the curve level.
@@
Despite a growing number of contributions, an essential question remains unanswered: Are merit-order curves forecasting models competitive with state-of-the-art price-centric EPF models in terms of clearing price prediction accuracy?
@@
Rigorously compare, in accordance with best practices established by lago_forecasting_2021, the clearing price forecasting performance of curve-based models, i.e. derived from merit-order curves predictions, with that of state-of-the-art price-based models, which directly model the clearing price.
@@
Our framework builds on ziel_electricity_2016 but eliminates the need for curves discretization by preserving their functional form. Specifically, we leverage functional principal component analysis to efficiently represent a pair of supply and demand curves in a vector space and employ regularized vector autoregressive models inspired by EPF literature for their prediction. We compare this functional approach with the original method of Ziel and Steinert and top-performing price-based point and probabilistic EPF models, on the Italian day-ahead market (GME), during the 2023-2024 period.
The remainder of this paper is structured as follows: section 2 reviews previous contributions, section 3 presents the knowledge gap addressed by this work, section 4 details the forecasting and testing methodology, section 5 describes the dataset, section 6 presents and discusses the empirical results, and section 7 provides concluding remarks.
2 A thematic review of prior work
Table 1 provides an overview of contributions on the topic. Rather than providing a linear and descriptive review, we organize it according to several key aspects that we believe are central to the merit-order curve forecasting problem.
2.1 Handling the infinite dimensionality
A central aspect of the problem regards the fact that merit-order curves are essentially infinite-dimensional functional objects, in the sense that they represent quantities that vary over a continuum. It is therefore necessary to consider either an appropriate statistical framework, e.g. functional data analysis (ramsay_functional_2005; ferraty_nonparametric_2006) or a finite-dimension representation of the curves to which classic multivariate time series methods can be applied.
Within the functional approach, several generalizations of the autoregressive (AR) models (box_time_1994) to the functional setting have been used: Functional AR (FAR) models (canale_constrained_2016; soloviova_efficient_2021; li_predicting_2025), functional SARMA (mestre_forecasting_2020) and nonparametric FAR (NPFAR) models (shah_forecasting_2020). Methodological reference on functional time series include bosq_linear_2000 and hormann_7_2012. The advantage of functional approaches, beyond elegantly tackling the problem from a mathematical standpoint, is that they avoid the information loss inherent in finite dimension representations. Their main drawback, however, lies in the limited theoretical development and the lack of efficient, easy-to-use, open-access implementations of the models proposed in the literature (aue_prediction_2015).
Aside of the functional framework, there are essentially three strategies to address the dimensionality problem: discretization, parametrization and dimensionality reduction. For all three, the overall workflow is similar: (i) Transform the curves into a finite-dimension vector form, (ii) Forecast this vector using multivariate methods, and (iii) Invert the transformation to recover the predicted curves.
Although any software ultimately relies on discrete representations, we refer to the discretization approach when the vector representation consists of curve evaluations over a grid of reasonable size (no more than a few dozen points), enabling multivariate modeling. The functional form is then recovered using a chosen interpolation method. This strategy was adopted in the seminal work of ziel_electricity_2016 and its extensions in kulakov_x-model_2020 and ghelasi_hierarchical_2024. Parametrization refers to methods that assume a parametric form (e.g. linear, logarithmic or polynomial) for the curves, with the finite-dimensional representation given by the fitted parameters; see for instance yildirim_supply_2023 and ciarreta_forecasting_2023. Finally, dimensionality reduction refers to approaches where the finite-dimension representation is obtained by means of a low-rank approximation of the data through standard matrix factorization techniques (guo_forecast_2021; tang_forecasting_2024), neural networks (sinha_demand_2025) or functional PCA (grigoletto_supply_2013; li_predicting_2025).
Functional PCA (FPCA) is a well-known, simple, and effective method for forecasting functional time series (hyndman_robust_2007; hyndman_forecasting_2009; shang_functional_2013; aue_prediction_2015). It is therefore a natural choice for a benchmark framework. Further arguments supporting this choice are provided in section 4.2.
Work Market Model typology Dim. reduction Curve form Monotonicity constraint Demand side Joint prediction Price forecasting Probabilistic forecasting grigoletto_supply_2013 GME VARX PCA Inverse Built-in ziel_electricity_2016 EPEX-DE VARX Discretization Inverse Built-in∗ ✓ ✓ ✓ ✓ canale_constrained_2016 GME (Gas market) Functional ✗ Original Built-in ✓ ✓ ✓ mestre_forecasting_2020 GME Functional ✗ Inverse Post-processing kulakov_x-model_2020 EPEX-DE VARX Discretization Inverse Built-in∗ ✓ ✓ ✓ shah_forecasting_2020 GME Functional ✗ Original ✗ ✓ ✓ ✓ guo_forecast_2021 MISO Neural Network PCA Original ✗ soloviova_efficient_2021 GME Functional ✗ Original Built-in∗ ✓ ✓ yildirim_supply_2023 EPIAS State-space Parametrization Inverse ✗ ciarreta_forecasting_2023 OMIE VARX Parametrization Original ✗ ✓ ✓ ✓ tang_forecasting_2024 ANEM Neural Network PCA, NMF, SDL Original ✗ vivo_prediction_2024 OMIE (Intraday market) Neural Network ✗ Inverse Post-processing ghelasi_hierarchical_2024 EPEX-DE VARX Discretization Inverse Built-in∗ ✓ ✓ li_clustering_2024 ANEM, OMIE, PJM k-NN ✗ Inverse Built-in ghelasi_data-driven_2025 EPEX-DE Fundamental ✗ Original Built-in ✓ sinha_demand_2025 GME Neural Network Autoencoder Original Post-processing ✓ li_predicting_2025 OMIE ARIMA, Functional FPCA Inverse Post-processing ✓ ✓
2.2 Modeling the original vs inverse curve
Being a monotone step-function, a merit-order curve expressing price as a function of quantity admits an inverse expressing quantity as a function of price which is also a monotone step function. These are just two different functional representations of the same data object, and one can easily switch from one representation to an other. Interestingly, a large number of contributions work with the inverse form – offering little discussion on the reasons behind this choice – though there is no clear consensus on what is the best representation strategy between the two forms. However, as we will see in section 4.1, the inverse form has a more natural mathematical definition and offers a convenient perspective on the main source of curves variability.
2.3 Addressing the monotonicity constraint
Merit-order curves are by definition monotone, increasing for supply and decreasing for demand and this must be ideally respected for predictions. To enforce this constraint, two strategies appear in the literature:
The first is to transform the curves to work in a space of unconstrained functions, usually achieved applying a composition of a logarithmic and differential operator on the smoothed curves (grigoletto_supply_2013; canale_constrained_2016), work with first-order differences over a discretization grid (ziel_electricity_2016; kulakov_x-model_2020; ghelasi_hierarchical_2024) or assuming a monotone parametric form (soloviova_efficient_2021; ciarreta_forecasting_2023; ghelasi_data-driven_2025). Note that works considering the two latter approaches do not exactly enforce the monotonicity constraint as the vector representation – either first-order differences or parametric coefficients – components sign is not constrained in the prediction process. It doesn’t seem however to be an issue in practice.
The second is to post-process the predictions, for instance by taking the predicted curve’s nearest neighbor in the training set (sinha_demand_2025; li_predicting_2025) or applying a correction re-establishing monotonicity in the predicted curve (mestre_forecasting_2020; vivo_prediction_2024).
In our case, rather than working in the constrained functional framework, although appealing from a mathematical point of view, we adopt the post-processing approach as the log-differential transformation was found to introduce excessive noise in the FPCA during preliminary experiments.
2.4 Dealing with the step-like characteristic
Merit-order curves are step functions whose step heights and widths vary across both the domain and time. Most studies, either explicitly or implicitly, work with continuous or smooth approximations of these curves – a practice we consider appropriate. Nevertheless, this approach has been criticized by several authors (vivo_prediction_2024; li_clustering_2024; li_predicting_2025; sinha_demand_2025), who therefore develop techniques that preserve the step-like form, particularly for point forecasts. However, this can lead to inconsistent or misguided modeling choices: For instance, li_clustering_2024 use a -NN algorithm to predict curves with to avoid the smoothing induced when averaging curves for , hence departing from the usual bias–variance trade-off considerations usually driving the choice of . Similarly, li_predicting_2025 use a functional basis of constant functions defined on equal-width, disjoint intervals, thereby imposing fixed step widths, which the data do not exhibit.
Beyond offering a more tractable form, working with smooth approximations is theoretically grounded. Indeed, as pointed out by grigoletto_supply_2013 and diquigiovanni_distribution-free_2024, step locations (the bid prices) are random; being real-valued, they can be modeled as realizations of a continuous random variable. Consequently the expected value of the curves process is a smooth function333Further arguments are provided in section 4.2. and smooth point forecasts – being conditional mean estimators – are therefore appropriate. This hypothesis is easily verified empirically by inspecting the smoothness of pointwise average curves in real merit-order curves datasets.
2.5 Joint prediction of supply and demand
Day-ahead power demand has historically shown little elasticity making the demand merit-order curve look more like a vertical line close to the load forecast than decreasing steps. This point has undoubtedly motivated academic researchers to focus primarily on the supply side (grigoletto_supply_2013; mestre_forecasting_2020; guo_forecast_2021; yildirim_supply_2023; tang_forecasting_2024; li_predicting_2025; ghelasi_data-driven_2025). However, first, this assumption does not hold for all markets (e.g. EPEX-DE, the German day-ahead market) and second, the increasing penetration of large-scale storage systems together with the growth of demand-side response is likely to increase day-ahead demand elasticity in the next years.
Although any methodology developed for supply curves is trivially applicable to demand curves by considering their opposite (which is a supply-like increasing function) the challenge is to account for the statistical dependence that may exist between the two objects. This point is crucial when the ultimate goal is price forecasting. Among the works that consider both supply and demand curves, only a few propose a method capable of making supply curves depend on past demand curves and vice-versa. A simple yet effective approach (which we adopt in this work) within dimensionality-reduction frameworks is to stack the supply and demand vector representations into a single vector representing the pair, to which multivariate forecasting models can then be applied (ziel_electricity_2016; ciarreta_forecasting_2023). This dependence structure is also incorporated by canale_constrained_2016 via a bivariate FAR model (a functional extension of a two-dimensional vector AR model) applied directly to the supply–demand function pair. A notable conclusion from their study is that modeling this dependence is indeed essential for achieving the best predictive performance.
2.6 Curves forecast evaluation and clearing price forecasting
A final crucial point regards the way merit-order curves forecasts are evaluated, since not all curves portions carry the same importance. Indeed, in a pay-as-clear auction, the bid price doesn’t matter so much as long as it is estimated to fall either below, above or around the market clearing price. As a result, as we move further from the clearing quantity, bid prices may be subject to an arbitrary variability which is irrelevant from an economic point of view. For instance, market participants with variable renewable energy assets may place bids either at -500€/MWh444The minimum bidding price in European markets. or 0€/MWh, while still being certain to be cleared if the market clearing price never falls below 10€/MWh. This variability is likely to be hard to capture and does not reflect variability in market outcomes.
A simple way to address this issue, adopted in several works, is to restrict the supply-demand curves pair to the narrowest domain (possibly with some margin) containing all historic intersections, i.e. clearing price-quantity points (mestre_forecasting_2020; shah_forecasting_2020; soloviova_efficient_2021). This is the approach we follow in this work. Since the conditional distribution of the clearing price exhibits smaller temporal variations than that of the clearing quantity – which has lower uncertainty and closely follows daily demand patterns, this approach is likely to be more effective at discarding low-significance bids when curves are represented in their inverse form. Another possible approach is to ”clip” bid prices within a conservative range of plausible clearing prices, as done by guo_forecast_2021 and tang_forecasting_2024. Alternative weighting-based evaluation methods have also been proposed, based on the distribution of bid prices (vivo_prediction_2024; li_clustering_2024) or clearing prices (tang_forecasting_2024). Their main drawback is that results tend to be highly sensitive to the chosen weighting scheme.
Finally, a desirable property of supply and demand curves forecasts is their ability to yield accurate clearing price predictions, and they must be evaluated accordingly. As shown in Table 1, most works addressing both supply and demand curves forecasting tested their ability to produce accurate price forecasts, but rarely following the EPF best practices established in lago_forecasting_2021.
3 Literature gap and contributions
3.1 Merit-order curves forecasting benchmark
As there are very few direct and fair comparisons between the different merit-order curves forecasting approaches proposed in the literature, and no shared guidelines or best practices for performance evaluation as in EPF, understanding which models work best remains difficult. Moreover, only a small number of papers provide indicative algorithms running times, preventing a clear assessment of their performance/efficiency trade-off.
ziel_electricity_2016 were the first to propose a method addressing all aspects of the problem. Their framework is simple, general enough to be applied across different markets, computationally efficient, and relatively easy-to-implement. Yet, aside from a few direct extensions (kulakov_x-model_2020; ghelasi_hierarchical_2024) and despite a large number of subsequent contributions, no alternative method has been conclusively shown to outperform the original approach of Ziel and Steinert. Their model therefore remains state-of-the-art and a natural – if not necessary – benchmark for comparison. We identified only one study, yildirim_supply_2023, that provides a direct comparison of forecast accuracy for supply curves. Their results suggest that their method and that of Ziel and Steinert perform similarly. A comparison is also reported in sinha_demand_2025, but the authors state that they employed a modified version of Ziel and Steinert’s model without providing further details. Additionally, we are somewhat skeptical of their results as all considered benchmarks – including Ziel and Steinert’s model – are found to perform far worse than a naive method.
This work aims to provide a general, simple, and computationally efficient benchmark framework for day-ahead supply and demand merit-order curves forecasting, from which both point and probabilistic clearing price predictions can be derived. We conduct a rigorous comparison with Ziel and Steinert’s framework and provide an open-access Python implementation555We also implement the model of ziel_electricity_2016, considering the multivariate models described in this paper – very similar to Ziel and Steinert’s original ones – within a shared software architecture. (github.com/gkchln/moc_forecast) to facilitate replication and application to other merit-order curves datasets.
3.2 Comparison with state-of-the-art EPF models for clearing price prediction
As mentioned earlier in the introduction, it remains unclear whether merit-order curves models are capable of yielding state-of-the-art point and probabilistic clearing price forecasting performance. Back in 2016, Ziel and Steinert showed that their model outperformed existing price-based approaches for point predictions, but this may no longer be the case, as of 2025.
Several works have validated the effectiveness of curve-based models for clearing price point forecasting by comparing them with price-based approaches, but the latter are often too simple and not representative of top-performing EPF methods. Moreover, the rigorous forecasts evaluation rules recommended by lago_forecasting_2021, such as daily recalibration, are seldom followed. Besides, none of the post-2021 contributions offer comparison with the benchmark models of Lago et al., despite their availability in an open-access software (epftoolbox Python package). shah_forecasting_2020 compare their curve-based price forecasts only against basic autoregressive models without exogenous covariates – even though such covariates are necessary for achieving top performance (weron_electricity_2014). A similar observation applies to soloviova_efficient_2021, ciarreta_forecasting_2023 and li_predicting_2025. ghelasi_data-driven_2025 consider an ”expert” autoregressive model including exogenous covariates, but do not incorporate regularized, parameter-rich autoregressive models with cross-hour dependence, despite evidence that these models consistently outperform expert ARX models (ziel_day-ahead_2018). For probabilistic forecasting, bootstrap-based Monte Carlo approaches similar to ours were proposed by ziel_electricity_2016 and shah_forecasting_2020, but they were not compared with state-of-the-art probabilistic EPF models.
A rigorous empirical comparison between our curve-based price forecasting approach and state-of-the-art price-based EPF models is conducted in this work. The comparison is restricted to regularized, parameter-rich autoregressive models (ziel_forecasting_2016; uniejewski_automated_2016; ziel_day-ahead_2018; lago_forecasting_2021) as our curves forecasting procedure relies on a direct multivariate extension applied to the curves vector representation. In these experiments, all models share a common regularized vector autoregression implementation in Python. For probabilistic forecasting, we compare with quantile regression averaging (nowotarski_recent_2018), conformal prediction (kath_conformal_2021) and isotonic distributional regression (henzi_isotonic_2021), using the open-access implementation of lipiecki_postprocessing_2024.

### v2

@@ Abstract
We introduce a general, simple, and computationally efficient framework for predicting day-ahead supply and demand merit-order curves, from which both point and probabilistic electricity price forecasts can be derived. We conduct a rigorous empirical comparison of price forecasting performance between the proposed curve-based model, i.e., derived from predicted merit-order curves, and state-of-the-art price-based models that directly forecast the clearing price, using data from the Italian day-ahead market over the 2023-2024 period. Our results show that the proposed curve-based approach significantly improves both point and probabilistic price forecasting accuracy relative to price-based approaches, with average gains of approximately 5%, and improvements of up to 10% during mid-day hours, when prices occasionally drop due to high renewable generation and low demand.
@@
energy forecasting , electricity , price forecasting , day-ahead market , merit-order curves , functional data analysis , multivariate time series , vector autoregression models , probability forecasting[mox]organization=MOX, Department of Mathematics, Politecnico di Milano,addressline=Piazza Leonardo da Vinci 32, city=Milano, postcode=20133, state=MI, country=Italy
@@ 1.1 Context
Because of this single-price property, a large body of research has focused on forecasting the day-ahead clearing price itself, a discipline known as electricity price forecasting (EPF). The consensus is that data-driven approaches based on historical prices tend to outperform fundamental (or structural) ones which try to model exact market mechanisms, especially for short-term hourly forecasts (Weron, 2014). In particular, a highly competitive day-ahead EPF benchmark based on regularized autoregressive models and feedforward neural networks was consolidated by Lago et al. (2021). This framework is still state-of-the-art and has been widely employed in recent studies (Olivares et al., 2023; Lipiecki et al., 2024; O’Connor et al., 2025; Cerasa and Zani, 2025).
@@
Nevertheless, the problem of forecasting the entire merit-order curves underlying price formation is gaining momentum in the academic community (Petropoulos et al., 2022). Main contributions on this topic can be found in Table 1. The seminal work of Ziel and Steinert (2016) demonstrated the capacity of ”data-driven fundamental” approaches based on historical auction data and their potential for generating competitive clearing price forecasts. Forecasting entire curves offers several advantages. For participants with market power, it enables precise quantification of how marginal changes in offers may influence the clearing price and allows realistic market simulations from which optimal bidding strategies can be derived (Pelagatti, 2013). For price-takers and other actors exposed to the day-ahead price, curve-based models offer a more comprehensive view of market results, enhances the interpretability of price predictions, and may improve the quantification of price predictions’ uncertainty222This latter aspect is investigated in this paper.. Additionally, complex nonlinear dynamics observed at the price level may become more tractable when modeled at the curve level.
Work Model typology Dim. reduction EPF Probabilistic EPF Pelagatti (2013) VAR FPCA Ziel and Steinert (2016) VAR Discretization ✓ ✓ Canale and Vantini (2016) Functional AR ✗ Mestre et al. (2020) Functional AR ✗ Kulakov (2020) VAR Discretization ✓ Shah and Lisi (2020) Functional AR ✗ ✓ ✓ Guo et al. (2021) Neural Network PCA Soloviova and Vargiolu (2021) Functional AR ✗ ✓ Yıldırım et al. (2023) State-space Parametrization Ciarreta et al. (2023) VAR Parametrization ✓ Tang et al. (2024) Neural Network PCA, NMF, SDL Vivó and Alonso (2024) Neural Network ✗ Ghelasi and Ziel (2024) VAR Discretization Li et al. (2024) k-NN ✗ Ghelasi and Ziel (2025) Fundamental ✗ ✓ Sinha and Lucheroni (2025) Neural Network Autoencoder Li et al. (2025) ARIMA, Functional AR FPCA ✓
1.2 Knowledge gap
Given the absence of a unified evaluation framework – as in EPF – and the scarcity of direct and fair comparisons between proposed approaches, identifying top-performing merit-order curves forecasting techniques remains challenging. In particular, aside from a few direct extensions (Kulakov, 2020; Ghelasi and Ziel, 2024) and despite a large number of subsequent contributions, no alternative method has been conclusively shown to outperform the original approach of Ziel and Steinert (2016). We identified only one study, Yıldırım et al. (2023), that provides a direct comparison, though limited to supply curve forecast accuracy. Their results suggest that their method and that of Ziel and Steinert perform similarly. A comparison of curve forecast performance is also reported in Sinha and Lucheroni (2025), although the authors state that they employ a modified version of Ziel and Steinert’s model without providing further details333Additionally, we are somewhat skeptical of their results as all considered benchmarks – including Ziel and Steinert’s model – are reported to perform far worse than a naive method.. As a result, Ziel and Steinert’s framework remains the de facto state-of-the-art, and whether it can be consistently improved remains an open question.
@@ 1.2 Knowledge gap
Beyond the problem of curve forecasting itself, an essential question remains unanswered: are merit-order curve forecasting models competitive with state-of-the-art price-centric EPF models in terms of clearing price prediction accuracy?
While Ziel and Steinert demonstrated ten years ago that their approach outperformed existing price-based models for point forecasting, this result may no longer hold in light of recent advances in EPF. Since then, several studies have assessed the effectiveness of curve-based approaches for clearing price prediction by comparing them with price-based methods; however, the latter are often overly simplistic and not representative of top-performing EPF models. In addition, the rigorous forecast evaluation practices recommended by Lago et al. (2021), such as daily recalibration, are seldom followed. Furthermore, none of the post-2021 contributions compare curve-based price forecasts against the benchmark models of Lago et al., despite their availability through an open-access software implementation (the epftoolbox Python package). For instance, Shah and Lisi (2020) compare their curve-based forecasts only against basic autoregressive models without exogenous covariates, even though variables such as load and renewable generation forecasts are known to be essential for achieving top performance (Weron, 2014). Similar limitations apply to Soloviova and Vargiolu (2021), Ciarreta et al. (2023), and Li et al. (2025). Ghelasi and Ziel (2025) consider an “expert” autoregressive model with exogenous covariates, but do not include regularized, parameter-rich autoregressive models with cross-hour dependence, despite evidence that such models consistently outperform expert ARX specifications (Ziel and Weron, 2018).
1.3 Contributions
@@
Rigorously compare the clearing price forecasting performance of the proposed curve-based models, i.e. derived from merit-order curves predictions, with that of state-of-the-art price-based models, which directly model the clearing price.
@@
Our framework builds on Ziel and Steinert (2016) but eliminates the need for curves discretization by preserving their functional form. Specifically, we leverage functional principal component analysis to efficiently represent a pair of supply and demand curves in a vector space and employ regularized vector autoregressive models inspired by EPF literature for their prediction. We compare this functional approach with the original method of Ziel and Steinert and top-performing price-based point and probabilistic EPF models, on the Italian day-ahead market (GME), during the 2023-2024 period.